Find the partial differential equation by eliminating a and b from the equation z=ax+by+a²+b²


1
Expert's answer
2021-06-16T12:27:30-0400
"z=ax+by+a^2+b^2"

Differentiating partially with respect to "x, y" we get


"\dfrac{\partial z}{\partial x}=p=a"

"\dfrac{\partial z}{\partial y}=q=b"

Then the partial differential equation is


"px+qy+p^2+q^2=z"